OpenStudy (anonymous):

Consider the function f(x) = 8 − 3x2. Find the value of f(x) when x = 3 and select the correct answer below. -19 -10 26 35

OpenStudy (anonymous):

the answer is 30

OpenStudy (anonymous):

but there is no 30 in the options

OpenStudy (anonymous):

okay let me try it again

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

its -10

OpenStudy (anonymous):

8-3(3)x2 8-9x2 8-18 -10

OpenStudy (phi):

is it \[ f(x) = 8 − 3x^2 \]? if so , the \(x^2\) means x times itself. in other words, you can write it this way \[ f(x) = 8 − 3\cdot x \cdot x \] to find f(3) you erase the x wherever you see it, and replace it with 3 then simplify (and remember "order of operations"... multiply before add)